write 3 different improper fractions that equal 4 2/3
Answer #1
Yay! You've got it! :-)
answered by: JOSHIKA
Answer #2
14/3; 28/6; 42/9

check those.
answered by: Marina
Answer #3
4 2/3 4 X 3 =12 (4 wholes into slices) 12 + 2 (2/3) =14/3 then multiply both numbers by 2 and you have various answers (do multiple times)
answered by: enrique
Answer #4
ok but how do you get those?
answered by: Roosevelt
Answer #5
got it. i just started mulitplying small numbers(like 3 & 4) by the numerator and denominator of 14/3(cause 14/3 is what you get when you turn mixed numders of 4 2/3 into improper fractions. so i got 42/9 & 56/12. correct?
Answer #6
4 2/3

Imagine a pizza cut into three equal pieces.

3/3 = 1
6/3 = 2
9/3 = 3
12/3 = 4
14/3 = 4 2/3

Then -- multiply the numerator and the denominator by the same number.
14/3 * 2/2 = 28/6
14/3 * 3/3 = 42/9
answered by: AnonymousII
Answer #7
three differnt fraction for 4 2/3
answered by: Emma